stylequeen:

The graph below shows a line segment AB: graph of line AB going through ordered pairs negative 1, 5, and 2, negative 4 Which of the following equations best represents the line segment AB? (5 points) y = −3x + 2 y = −2x + 1 y = −3x − 2 y = −2x − 1

stylequeen:

can someone help me with this problem

supie:

Sure!

sandra17:

y= 3x+2

stylequeen:

can u plz explain it

sandra17:

= (y2 - y1)/(x2 - x1) (y - 5)/(x - (-1)) = (-4 - 5)/(2 - (-1)) (y - 5)/(x + 1) = -9/(2 + 1) = -9/3 = -3 y - 5 = -3(x + 1) = -3x - 3 y = -3x - 3 + 5 y = -3x + 2

stylequeen:

ok thank u have a good day

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sandra17 = (y2 - y1)/(x2 - x1) (y - 5)/(x - (-1)) = (-4 - 5)/(2 - (-1)) (y - 5)/(x + 1) = -9/(2 + 1) = -9/3 = -3 y - 5 = -3(x + 1) = -3x - 3 y = -3x - 3 + 5 y = -3x + 2 \(\color{#0cbb34}{\text{End of Quote}}\) Don't copy from other websites, it's against the rules.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) @sandra17 = (y2 - y1)/(x2 - x1) (y - 5)/(x - (-1)) = (-4 - 5)/(2 - (-1)) (y - 5)/(x + 1) = -9/(2 + 1) = -9/3 = -3 y - 5 = -3(x + 1) = -3x - 3 y = -3x - 3 + 5 y = -3x + 2 \(\color{#0cbb34}{\text{End of Quote}}\) Don't copy from other websites, it's against the rules. \(\color{#0cbb34}{\text{End of Quote}}\) Especially copying and taking it as your own work.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) supie \(\color{#0cbb34}{\text{Originally Posted by}}\) supie \(\color{#0cbb34}{\text{Originally Posted by}}\) sandra17 = (y2 - y1)/(x2 - x1) (y - 5)/(x - (-1)) = (-4 - 5)/(2 - (-1)) (y - 5)/(x + 1) = -9/(2 + 1) = -9/3 = -3 y - 5 = -3(x + 1) = -3x - 3 y = -3x - 3 + 5 y = -3x + 2 \(\color{#0cbb34}{\text{End of Quote}}\) Don't copy from other websites, it's against the rules. \(\color{#0cbb34}{\text{End of Quote}}\) Especially copying and taking it as your own work. \(\color{#0cbb34}{\text{End of Quote}}\) Also, especially if its from another HomeWork site.

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie \(\color{#0cbb34}{\text{Originally Posted by}}\) supie \(\color{#0cbb34}{\text{Originally Posted by}}\) supie \(\color{#0cbb34}{\text{Originally Posted by}}\) sandra17 = (y2 - y1)/(x2 - x1) (y - 5)/(x - (-1)) = (-4 - 5)/(2 - (-1)) (y - 5)/(x + 1) = -9/(2 + 1) = -9/3 = -3 y - 5 = -3(x + 1) = -3x - 3 y = -3x - 3 + 5 y = -3x + 2 \(\color{#0cbb34}{\text{End of Quote}}\) Don't copy from other websites, it's against the rules. \(\color{#0cbb34}{\text{End of Quote}}\) Especially copying and taking it as your own work. \(\color{#0cbb34}{\text{End of Quote}}\) Also, especially if its from another HomeWork site. \(\color{#0cbb34}{\text{End of Quote}}\) Thats plagiarism, and plagiarism is academic dishonesty, so don't copy from any other websites...